# CryoBench: Diverse and challenging datasets for the heterogeneity problem in cryo-EM

### v1

@@ Abstract
Cryo-electron microscopy (cryo-EM) is a powerful technique for determining high-resolution 3D biomolecular structures from imaging data. As this technique can capture dynamic biomolecular complexes, 3D reconstruction methods are increasingly being developed to resolve this intrinsic structural heterogeneity. However, the absence of standardized benchmarks with ground truth structures and validation metrics limits the advancement of the field. Here, we propose CryoBench, a suite of datasets, metrics, and performance benchmarks for heterogeneous reconstruction in cryo-EM. We propose five datasets representing different sources of heterogeneity and degrees of difficulty. These include conformational heterogeneity generated from simple motions and random configurations of antibody complexes and from tens of thousands of structures sampled from a molecular dynamics simulation. We also design datasets containing compositional heterogeneity from mixtures of ribosome assembly states and 100 common complexes present in cells. We then perform a comprehensive analysis of state-of-the-art heterogeneous reconstruction tools including neural and non-neural methods and their sensitivity to noise, and propose new metrics for quantitative comparison of methods. We hope that this benchmark will be a foundational resource for analyzing existing methods and new algorithmic development in both the cryo-EM and machine learning communities.
@@ 1 Introduction
Figure 1: Overview of CryoBench. (a) In cryo-EM, each image $X_{i}$ captures a molecule $V_{i}$ projected at an unknown pose $\phi_{i}$. A latent variable $z_{i}$ models the conformational space $\mathcal{V}$ that describes the heterogeneity between molecules $\{V_{i}\}$. (b) CryoBench includes 5 synthetic datasets of varying difficulty, characterized by heterogeneity arising from conformational (i.e. shape) or compositional (i.e. identity) changes. (c) Methods can be grouped into using either a continuous latent variable $z$ or discrete latent variable $\pi$ for heterogeneity. Hidden variables assumed to be known are shown in gray. Volumes are represented as a neural field (NF), voxel array (VA), neural volume (NV), or tetrahedral mesh (TM). Generative models are colored blue for nonlinear neural methods; orange for linear generative models, pink for mixture models; and green for density-preserving motion models. (d) A variety of metrics are used to assess both latent inference and volume reconstruction quality.
Previous methods development has relied on various types of benchmarks. These include: a) real datasets known to have sensible conformal trajectories, allowing experts to perform qualitative benchmarking by visually assessing reconstructed volumes, b) datasets of synthetic blob-like volumes used to demonstrate that models’ conformational latent spaces can represent simple 1D and 2D motions as parameterized motions, and c) pseudo-real motions, where real molecular structures are used, but simple motions such as rotations are applied to the subregions of the structure to generate a continuum of conformations. The lack of realistic and common benchmarks poses challenges in training models whose performance can generalize well across different datasets or in comparing existing heterogeneous reconstruction methods without relying on expert intuition. Moreover, because each method has been applied to different datasets designed to showcase different types of heterogeneity, practitioners cannot compare and determine the most promising method for their application. Finally, since these methods are mostly tested on real datasets with no ground truth, it’s hard to assess whether a given method produces accurate results on a new dataset, or how much to trust its scientific conclusions.
@@ 2 Background and Related Work
Past Benchmarks. While standard benchmarks in this field have yet to be established, there are some consistent datasets that have been used. For example, an experimental dataset that has been used for compositional heterogeneity is the different assembly states of the large subunit of a bacterial ribosome (EMPIAR-10076) [26], and a commonly used real dataset for conformational heterogeneity is the pre-catalytic spliceosome (EMPIAR-10180) [27]. However, it is also common to generate simulated cryo-EM datasets for which a ground truth is known [28, 6, 29, 30], but so far these are generated ad hoc for a research project, and not taken from a standard benchmark dataset for which the performance of other methods is known. There have been attempts at benchmark studies for heterogeneity where many methods are applied to a single dataset [31, 32], but this approach does not address performance across different types of heterogeneity such as conformational motions or compositional changes. Ideally, a common set of diverse benchmark datasets can be used by the community to compare different methods. Here, we focus on synthetic but challenging datasets where ground truth information is available for all latent variables, allowing for rigorous, quantitative benchmarking.
@@
Metrics. In addition to benchmark datasets, metrics with which to assess heterogeneity methods themselves are also lacking. Metrics for assessing homogeneous 3D reconstruction methods, where only a single volume is achieved from a cryo-EM image stack, are more mature and often incorporate the Fourier shell correlation (FSC) to judge resolution  [33, 34], and a gold standard approach for computing [35] has been widely adopted by the community (though this has its own degree of controversy [36]). However, when it comes to evaluating heterogeneous reconstructions, metrics become much less straightforward, and the standard FSC-based approach is flawed as it provides a global assessment of resolution and is typically performed on independent half-sets and thus is only a self-consistency measure. Recent work has introduced new metrics either to improve priors for variational autoencoders or to disentangle latent embeddings based on ground truth [37, 38, 39]. However, analysis metrics that can compare different methods remain challenging. Here, we choose to use metrics that depend on knowing ground truth from synthetic datasets, which either use (a) the FSC —but with nuances appropriate to for heterogeneity—or (b) applying local neighborhood comparisons and clustering accuracy on latent representations  [40, 41].
@@
Figure 2: IgG-1D results. (a) Dataset design with 100 ground truth (G.T.) conformations producing a circular motion. (b) A representative volume for each method with the G.T. volume. (c) Latent embeddings visualized by UMAP and colored by the G.T. dihedral angle parameterizing the circular motion. (d) Embedding comparison metrics. (e) Per-conformation FSC. Each curve shows the average FSC curve across all conformations with error bars indicating the standard deviation. Colors in (b), (d), and (e) correspond to methods shown in the legend. Additional results shown in Figure S7 and S13.
@@ 3.1 Image Formation Model
where $C_{i}$ is the Contrast Transfer Function (CTF), $P_{\phi}$ is a slice operator corresponding to an tomographic projection of the volume $V$ according to the pose $\phi=(R,\mathbf{t})\in\text{SO}(3)\times\mathbb{R}^{2}$, and $\eta_{i}$ is additive white Gaussian noise. Images are sampled on a $D\times D$ grid with $D=128$ by default unless otherwise specified. Additional details for each dataset are below with full dataset generation details given in the Appendix A.
Figure 3: IgG-1D with noise.
(a) Per-Conformation FSC for each method at SNR levels of 0.01, 0.005, 0.001 and example cryo-EM images. (b) UMAP visualizations colored by G.T. dihedral conformations of two representative methods. Additional results shown in Figure S11S16, and S17.
@@ 3.2 Conformational Heterogeneity
IgG-1D. Starting from an atomic model of the human immunoglobulin G (IgG) antibody (PDB: 1HZH), conformational heterogeneity is produced by rotating a dihedral angle connecting one of the fragment antibody (Fab) domains (Fig. 2a), simulating a simple one-dimensional continuous circular motion. This process yields 100 atomic models approximating the continuous dihedral rotation (360 degrees, 3.6-degree intervals). We simulate 1,000 images for each conformation, apply CTF and add noise at a signal-to-noise (SNR) ratio of 0.01 to produce a dataset of 100k images.
IgG-RL. Many protein complexes, including IgG, possess relatively rigid domains connected by a disordered peptide linker (e.g. exemplifying “beads on a string”). To model this more realistic and complex motion and provide a challenging case of heterogeneity, we generate random conformations for the linker connecting the Fab to the rest of the IgG complex (Fig. 4b). We identified a sequence of 5 residues as the linker and generated random realizations of its structure by sampling the backbone dihedral angles according to the Ramachadran distributions of disordered peptides from Towse et al. [42], using rejection sampling to eliminate structures with steric clashes. We generate 100 such atomic models and simulate 1k projections per conformation, apply CTF and add noise at an SNR of 0.01 to produce a dataset of 100k images.
Spike-MD. As a challenging example of detailed motions, we use a long-timescale molecular dynamics simulation as a source of ground truth models for dataset construction. Specifically, we use a simulation of the activating SARS-CoV-2 spike protein from Wieczór et al. [43]. Images for Spike-MD are generated at a higher resolution ($D=256$) to assess the ability of methods to capture high-resolution motions. We generate a dataset of 100k projection images at an SNR of 0.1.
IgG-1D-noisier and -noisiest. One of the most salient characteristics of cryo-EM images is the high degree of noise. As a test of the robustness of each method to the amount of noise, we also create versions of the IgG-1D dataset at SNR 0.005 for IgG-1D-noisier and 0.001 for IgG-1D-noisiest.
@@ 3.3 Compositional Heterogeneity
Figure 4: IgG-RL results. (a) Dataset design. The dihedral angles of a linking peptide are chosen to randomly orient one arm of the IgG molecule in 100 unique conformations. (b) The distribution of the 100 structures across coordinates characterized by the angle and distance between the rotating arm and the rigid arms. The UMAP plots of RECOVAR and OPUS-DSD latent spaces colored by the distance G.T. coordinate. (c) A representative volume for each method with the G.T. volume. (d) Embedding Comparison metrics. (e) Per-Conformation FSC curves. Each curve shows the average FSC curve across all conformations with error bars indicating the standard deviation. Colors in (d), (e), and (f) correspond to methods shown in the legend. Additional results shown in Figure S12 and S14.
Tomotwin-100. While cryo-EM is typically performed on purified samples or simple mixtures, in principle, the technique can be used to image complex mixtures e.g. from cellular lysates or in situ samples from cryo-ET. To create a challenging dataset of modeling compositional heterogeneity of this scale, we use atomic models from the curated set of cellular complexes in Rice et al. [45]. Here we use 100 out of the original set of 120 after excluding the 15 smallest and 5 largest complexes. Figure 7a shows 10 out of 100 ground truth volumes colored by size (i.e. molecular weight).
@@ 4 Evaluation Framework
As part of CryoBench, we consider methods for heterogeneous reconstruction with fixed (ground truth) poses, and the more challenging task of ab initio reconstruction where no input poses are provided. We evaluate seven fixed pose state-of-the-art methods and three ab initio variants. The fixed pose methods include 3D Classification (3D Class) in cryoSPARC [46], 3DVA [8], 3DFlex [9], CryoDRGN [7], DRGN-AI-fixed [47], Opus-DSD [18], and RECOVAR [19]. For ab initio methods, we use multiclass ab initio reconstruction in cryoSPARC (3D Class abinit) [46], CryoDRGN2 [48], and DRGN-AI [47]. An overview of these methods and training details can be found in Appendix B.
Figure 5: Spike-MD results. (a) Dataset design. 46,789 structures were sampled from a MD simulation of the SARS-COV-2 spike protein, including opening of the receptor binding domain (RBD, shown in red). The motion in the MD simulation can be described with two collective variables (CV), corresponding to opening and twisting of the RBD. (b) The population density of molecular states projected onto these CVs. (c) A representative volume for each method with the G.T. volume. (d) Latent embeddings visualized by UMAP and colored by the first and second CV. Additional results shown in Figure S8,  S15, and  S22.
@@ 4.1 Analysis and Metrics
For quantitative evaluation, we propose three metrics for comparison of embeddings: 1) Neighborhood Similarity, 2) Information imbalance, and 3) Adjusted Rand Index (ARI). To assess reconstructed volumes, we propose “Per-Conformation Fourier Shell Correlation (FSC)” as a metric that jointly evaluates image conformation estimation and reconstruction quality.
Figure 6: Ribosembly results. (a) 16 G.T. Ribosome assembly states. (b) A representative volume for each method with the G.T. volume. (c) Latent embeddings visualized by UMAP and colored by the 16 G.T. states. (d) Embedding Comparison metrics. (e) Per-Conformation FSC plots. Each curve shows the average FSC curve across all conformations with error bars indicating the standard deviation. Colors in (b), (d), and (e) correspond to methods shown in the legend. Additional results shown in Figure S18.
@@ 4.1.1 Embedding Comparisons
Neighborhood Similarity. To quantify the similarity between local neighborhoods in the learned embedding spaces and a ground truth heterogeneity embedding, we first use a generalization of neighborhood similarity [40]. We quantify the percentage of matching neighbors (pMN) with respect to the ground truth that are found within a neighborhood radius $k$,
@@
Clustering Accuracy. To assess each method’s ability to separate compositional states, we first use $k$-means clustering on the latent embeddings with $k$ equal to the number of ground truth states ($k=16$ for Ribosembly and $k=100$ for Tomotwin-100). We then compare these cluster assignments to the true structural labels for each particle using two metrics commonly employed for clustering consistency, the Adjusted Rand Index (ARI) and Adjust Mutual Information (AMI) [49] [50]. We note that this metric is sensitive to the performance of the clustering, and might differ with alternative algorithms to $k$-means.
@@ 4.1.2 Volume Metrics
We use Fourier Shell Correlation (FSC), a standard volume comparison metric in cryo-EM, to evaluate reconstruction quality against ground truth structures. Typically, FSC compares two independent reconstructions from separate dataset halves, with resolution determined at an FSC cutoff of 0.143 [35]. Here, we compare reconstructed volumes against ground truth volumes and compute the Area Under the FSC Curve (AUC-FSC) as a summary statistic. We show that compared to an FSC cutoff, the AUC is more sensitive to structural differences amongst the ground truth volumes (Appendix C.1.2, Fig S4).
For heterogeneity analysis, we propose the Per-Conformation FSC to jointly assess conformation estimation and reconstruction quality. For each ground truth conformation $V$, we compute a corresponding latent coordinate by averaging all latent embeddings $\bar{z}$ of images generated by $V$. To remain on the data manifold, we identify the closest latent coordinate $z^{*}$ to $\bar{z}$. Using this coordinate, we reconstruct the associated volume and evaluate its AUC-FSC against the ground truth volume. We also compute two additional settings that use other methods for choosing representative volumes for volume metrics: Sample FSC and Per-image FSC. See Appendix C.2 and Figure S9-S10 for more details.

### v2

@@ Abstract
Cryo-electron microscopy (cryo-EM) is a powerful technique for determining high-resolution 3D biomolecular structures from imaging data. Its unique ability to capture structural variability has spurred the development of heterogeneous reconstruction algorithms that can infer distributions of 3D structures from noisy, unlabeled imaging data. Despite the growing number of advanced methods, progress in the field is hindered by the lack of standardized benchmarks with ground truth information and reliable validation metrics. Here, we introduce CryoBench, a suite of datasets, metrics, and benchmarks for heterogeneous reconstruction in cryo-EM. CryoBench includes five datasets representing different sources of heterogeneity and degrees of difficulty. These include conformational heterogeneity generated from designed motions of antibody complexes or sampled from a molecular dynamics simulation, as well as compositional heterogeneity from mixtures of ribosome assembly states or 100 common complexes present in cells. We then analyze state-of-the-art heterogeneous reconstruction tools, including neural and non-neural methods, assess their sensitivity to noise, and propose new metrics for quantitative evaluation. We hope that CryoBench will be a foundational resource for accelerating algorithmic development and evaluation in the cryo-EM and machine learning communities. Project page: https://cryobench.cs.princeton.edu.
@@ 1 Introduction
Figure 1: Overview of CryoBench. a) Image formation model. In cryo-EM, each image $X_{i}$ captures a molecule $V_{i}$ projected at an unknown pose $\phi_{i}$. A latent variable $z_{i}$ models the conformational space $\mathcal{V}$ that describes the heterogeneity among the set of molecules $\{V_{i}\}$. b) Datasets. CryoBench includes 5 synthetic datasets of varying difficulty, characterized by heterogeneity arising from conformational (i.e. shape) or compositional (i.e. identity) changes. c) Methods. Methods can be grouped into using either a continuous latent variable $z$ or discrete latent variable $\pi$ for modeling heterogeneity. Hidden variables assumed to be known are shown in gray. Volumes are represented as a neural field (NF), voxel array (VA), neural volume (NV), or tetrahedral mesh (TM). Generative models are colored blue for nonlinear neural methods; orange for linear generative models, pink for mixture models; and green for density-preserving motion models. d) Metrics. Summary of metrics used to assess both latent inference and volume reconstruction quality.
Previous methods development has relied on various types of datasets for benchmarking and validation. These include: a) real datasets known to have sensible conformal trajectories, allowing experts to perform qualitative benchmarking by visually assessing reconstructed volumes, b) datasets of synthetic blob-like volumes used to demonstrate that models’ conformational latent spaces can represent simple 1D and 2D motions as parameterized motions, and c) pseudo-real motions, where real molecular structures are used, but simple motions such as rotations are applied to the subregions of the structure to generate a continuum of conformations. The lack of realistic and common benchmarks poses challenges in training models whose performance can generalize well across different datasets or in comparing existing heterogeneous reconstruction methods without relying on expert intuition. Moreover, because each method has been applied to different datasets designed to showcase different types of heterogeneity, practitioners cannot compare and determine the most promising method for their application. Finally, since these methods are mostly tested on real datasets with no ground truth, it’s hard to assess whether a given method produces accurate results on a new dataset, or how much to trust its scientific conclusions.
@@ 2 Background and Related Work
Past Benchmarks. While standard benchmarks in this field have yet to be established, there are some consistent datasets that have been used. For example, an experimental dataset that has been used for compositional heterogeneity is the different assembly states of the large subunit of a bacterial ribosome (EMPIAR-10076) [26], and a commonly used real dataset for conformational heterogeneity is the pre-catalytic spliceosome (EMPIAR-10180) [27]. However, it is also common to generate simulated cryo-EM datasets for which a ground truth is known [28, 6, 29, 30], but so far these are generated ad hoc for a given study, and not taken from a standard benchmark dataset for which the performance of other methods is known. There have been attempts at benchmark studies for heterogeneity where many methods are applied to a single dataset [31, 32], but this approach does not address performance across different types of heterogeneity such as conformational motions or compositional changes. Ideally, a common set of diverse benchmark datasets can be used by the community to compare different methods. Here, we focus on synthetic but challenging datasets where ground truth information is available for all latent variables, allowing for rigorous, quantitative benchmarking.
@@
Metrics. In addition to benchmark datasets, metrics with which to assess heterogeneity methods themselves are also lacking. Metrics for assessing homogeneous 3D reconstruction methods, where only a single volume is achieved from a cryo-EM image stack, are more mature and often incorporate the Fourier shell correlation (FSC) to judge resolution  [33, 34], and a gold standard approach for computing FSC [35] has been widely adopted by the community (though this has its own degree of controversy [36]). However, when it comes to evaluating heterogeneous reconstructions, metrics become much less straightforward, and the standard FSC-based approach is flawed as it provides a global assessment of resolution and is typically performed on independent half-sets and thus is only a self-consistency measure. Recent work has introduced new metrics either to improve priors for variational autoencoders or to disentangle latent embeddings based on ground truth [37, 38, 39]. However, analysis metrics that can compare different methods remain challenging. Here, we choose to use metrics that depend on knowing ground truth from synthetic datasets, which either use (a) a distributional volume reconstruction quality metric based on the FSC or (b) applying local neighborhood comparisons and clustering accuracy on latent representations [40, 41].
@@
Figure 2: IgG-1D results. a) Dataset design. Conformational heterogeneity of an IgG antibody complex produced from a simple, 1D continuous circular motion. b) Representative reconstructed and ground truth (G.T.) volumes. c) Latent embeddings visualized by UMAP and colored by the G.T. dihedral angle parameterizing the circular motion. Discrete class assignments are plotted by G.T. dihedral angles. d) Latent embedding analysis by neighborhood similarity and information imbalance. e) Per-Image FSC curves. Each curve shows the average FSC curve across all conformations with error bars indicating the standard deviation. Colors in b), d), and e) correspond to methods shown in the legend. Additional results shown in Figure S15.
@@ 3.1 Image Formation Model
where $C_{i}$ is the Contrast Transfer Function (CTF), $P_{\phi}$ is a slice operator corresponding to an tomographic projection of the volume $V$ according to the pose $\phi=(R,\mathbf{t})\in\text{SO}(3)\times\mathbb{R}^{2}$, and $\eta_{i}$ is additive white Gaussian noise. Images are sampled on a $D\times D$ grid with $D=128$ by default unless otherwise specified. Additional details for each dataset are below with full dataset generation details given in Appendix A.
@@ 3.2 Conformational Heterogeneity
IgG-1D. We use an atomic model of the human immunoglobulin G (IgG) antibody complex (PDB: 1HZH). Conformational heterogeneity is produced by rotating a dihedral angle connecting one of the fragment antibody (Fab) domains (Fig. 2a), simulating a simple one-dimensional continuous circular motion. This process yields 100 atomic models approximating the continuous dihedral rotation (at 3.6-degree intervals). We simulate 1,000 projection images for each conformation, apply CTF and add noise at a signal-to-noise (SNR) ratio of 0.01 to produce a dataset of 100k images.
Figure 3: IgG-1D with noise. a) Per-Image FSC for each method at different noise levels. Markers correspond to the legend in Figure 2. b) Example cryo-EM images for different noise levels and latent embeddings visualized by UMAP for CryoDRGN-AI. Additional results shown in Figure S13S18, and S19.
IgG-1D-noisier and -noisiest. As one of the most salient characteristics of cryo-EM images is the high degree of noise, we also create versions of the IgG-1D dataset at SNR 0.005 for IgG-1D-noisier and 0.001 for IgG-1D-noisiest to test the robustness of each method to the amount of noise.
IgG-RL. Many protein complexes, including IgG, possess relatively rigid domains connected by a disordered peptide linker (e.g. exemplifying “beads on a string”). To model this more realistic and complex motion and provide a challenging case of heterogeneity, we generate random conformations for the linker connecting the Fab to the rest of the IgG complex (Fig. 4b). We identified a sequence of 5 residues as the linker and generated random realizations of its structure by sampling the backbone dihedral angles according to the Ramachadran distributions of disordered peptides [42], using rejection sampling to eliminate structures with steric clashes. We generate 100 such atomic models and simulate 1k projections per conformation, apply CTF and add noise at an SNR of 0.01 to produce a dataset of 100k images.
Spike-MD. For a challenging dataset containing detailed motions, we use a long-timescale molecular dynamics simulation as a source of ground truth models for dataset construction. Specifically, we use a simulation of the activating SARS-CoV-2 spike protein from Wieczór et al. [43]. Images for Spike-MD are generated at a higher resolution ($D=256$) to assess the ability of methods to capture high-resolution motions. We generate a dataset of 100k projection images at an SNR of 0.1.
@@ 3.3 Compositional Heterogeneity
Figure 4: IgG-RL results. a) Dataset design. Conformational heterogeneity is produced by sampling 100 configurations of a peptide linker, randomly orienting the FAb domain in the IgG antibody complex. b) Representative reconstructed and ground truth (G.T.) volumes. c) The UMAP plots of RECOVAR and OPUS-DSD latent spaces colored by the distance between the FAb and the Fc domain in the G.T. volumes. d) Latent embedding analysis by neighborhood similarity and information imbalance. e) Per-Image FSC curves. Each curve shows the average FSC curve across all conformations with error bars indicating the standard deviation. Colors in (d), (e), and (f) correspond to methods shown in the legend. Additional results shown in Figure S14 and S16.
Tomotwin-100. While cryo-EM is typically performed on purified samples or simple mixtures, in principle, the technique can be used to image complex mixtures e.g. from cellular lysates or in situ samples from cryo-ET. To create a challenging dataset of modeling compositional heterogeneity of this scale, we use atomic models from the curated set of cellular complexes in Rice et al. [45]. Here we use 100 out of the original set of 120 after excluding the 15 smallest and 5 largest complexes. Figure 7a shows 10 out of 100 ground truth volumes colored by size.
@@ 4 Evaluation Framework
As part of CryoBench, we consider methods for heterogeneous reconstruction with fixed (ground truth) poses, and the more challenging task of ab initio reconstruction where no input poses are provided. We evaluate seven fixed pose state-of-the-art methods and three ab initio variants. The fixed pose methods include 3D Classification (3D Class) in cryoSPARC [46], 3DVA [8], 3DFlex [9], CryoDRGN [7], CryoDRGN-AI-fixed [47], Opus-DSD [18], and RECOVAR [19]. For ab initio methods, we use multiclass ab initio reconstruction in cryoSPARC (3D Class abinit) [46], CryoDRGN2 [48], and CryoDRGN-AI [47]. An overview of these methods and training details can be found in Appendix B.
Figure 5: Spike-MD results. a) Dataset design. 46,789 structures were sampled from a MD simulation of the SARS-COV-2 spike protein, including opening of the receptor binding domain (RBD, shown in red). The motion in the MD simulation can be described with two collective variables (CV), corresponding to opening and twisting of the RBD. b) The population density of molecular states projected onto these CVs. c) Representative reconstructed and ground truth (G.T.) volumes. d) Latent embeddings visualized by UMAP and colored by the first and second CV. Additional results shown in Figure S9,  S17, and  S24.
@@ 4.1 Analysis and Metrics
For quantitative evaluation, we propose three metrics for comparison of embeddings: 1) Neighborhood Similarity, 2) Information imbalance, and 3) Adjusted Rand Index (ARI). To assess reconstructed volumes, we use “Per-Image Fourier Shell Correlation (FSC)” as a metric that jointly evaluates image conformation estimation and reconstruction quality [6]. Moreover, we compute the pose error for ab initio methods in Appendix C.7.
Figure 6: Ribosembly results. a) Dataset design. 16 ground truth (G.T.) ribosome assembly states. b) Representative reconstructed and G.T. volumes. c) Latent embeddings visualized by UMAP and colored by the 16 G.T. states. d) Latent embedding analysis by neighborhood similarity and information imbalance. e) Per-Image FSC curves. Each curve shows the average FSC curve across all conformations with error bars indicating the standard deviation. Colors in (b), (d), and (e) correspond to methods shown in the legend. Additional results shown in Figure S20.
@@ 4.1.1 Embedding Comparisons
Neighborhood Similarity. To quantify the similarity between local neighborhoods of the learned embedding spaces and a ground truth heterogeneity embedding, we first use a generalization of neighborhood similarity [40]. We quantify the percentage of matching neighbors (pMN) with respect to the ground truth that are found within a neighborhood radius $k$,
@@
Clustering Accuracy. To assess each method’s ability to separate compositional states, we first use $k$-means clustering on the latent embeddings with $k$ equal to the number of ground truth states ($k=16$ for Ribosembly and $k=100$ for Tomotwin-100). We then compare these cluster assignments to the true structural labels for each particle using the Adjusted Rand Index (ARI) and Adjust Mutual Information (AMI) [49] [50]. We note that this metric is sensitive to the performance of the clustering, and might differ with alternative algorithms to $k$-means.
@@ 4.1.2 Volume Metrics
We use Fourier Shell Correlation (FSC), a standard volume comparison metric in cryo-EM, to evaluate reconstruction quality against ground truth structures. Typically, FSC compares two independent reconstructions from separate dataset halves, with resolution determined at an FSC cutoff of 0.143 [35]. Here, we compare reconstructed volumes against ground truth volumes and compute the Area Under the FSC Curve ($\text{FSC}_{\text{AUC}}$) as a summary statistic. We show that compared to an FSC cutoff, the AUC is more sensitive to structural differences amongst the ground truth volumes (Appendix C.1.2, Fig S5).
For heterogeneity analysis, we use the Per-Image FSC [6] as a distributional metric that jointly assesses conformation estimation and reconstruction quality. Here, we use a set of $N$ images from the dataset (either sampled at random or uniformly for each conformation). For each image, we reconstruct an associated volume at its inferred latent coordinate. We then compute the $\text{FSC}_{\text{AUC}}$ between the reconstructed volume and the ground truth volume for the image. We also investigated two alternative methods for sampling representative volumes to compute volume metrics: Sample FSC and Per-Conformation FSC. See Appendix C.2 and Figure S8-S11 for more details.